# Consumption Partial Insurance

### v3

@@ 5.4 Sensitivity to age, wealth, and education
We next consider two subsamples formed on the basis of household wealth. We define a household as ‘low wealth’ (‘high wealth’) if its average wealth over its entire observation window is less (more) than median real wealth in the sample over 19991999199919992019201920192019. Wealth is the sum of the value of housing, vehicles, business, investment, pension, savings, and other assets, net of mortgage and medical, student, and credit card debt. Table 1 provides summary statistics for the wealth subsamples. Average wealth in the first group is \$69,731 in 2018 prices, while the second group has 14 times more wealth, an average of \$950,823. Income and earnings are twice as high in the second group; consumption is only 30% higher.
Table 6, columns 1-2, presents the results in the quadratic model. [^44] Appendix table C.8, columns 1-4, presents the parameter estimates in the linear model. Three points emerge. First, $\phi^{(1)}$ is more than twice as large among the low vis-à-vis high wealth households ($0.212$ versus $0.091$). As $\phi^{(1)}$ measures the transmission of the average permanent shock, low wealth households exhibit a larger pass-through of average shocks, i.e., they are less able to insure against them, compared to high wealth households. Second, $\phi^{(2)}<0$ for both groups, albeit statistically insignificant. With the caveat of insignificance, the magnitude of $\phi^{(2)}$ is bigger among the wealthy. Large negative permanent shocks thus impact consumption of the wealthy more severely than they do among the less wealthy. As in the case of older households previously, there is more room for consumption to fall sharply among the wealthier. Third, $\psi^{(2)}$, the transmission parameter of the typical transitory shock away from the mean, is negative, large, and statistically significant among the less wealthy. Transitory shocks thus have an asymmetric pass-through among lower wealth households, with large bad shocks reducing consumption substantially more than good shocks increase it. So while average transitory shocks are fully insured among low wealth households, larger negative transitory shocks are not: they reduce consumption quite substantially as low wealth households lack the self-insurance means to smooth out even transitory fluctuations.
@@
Subsample: Low wealth High wealth No college Some college
@@
Notes: The table presents the estimates of the parameters of the quadratic consumption function, allowing them to vary across wealth and education groups (linear specification estimates appear in appendix table C.8). Low (high) wealth is defined on the basis of household wealth being less (more) than median real wealth in the sample over 19991999199919992019201920192019. No versus some college is defined on the basis of the highest level of education attained by the male spouse. All columns use consumption data internally available in the PSID. Estimation is via equally weighted GMM; block bootstrap standard errors are in parentheses.
We finally turn to the role of education. We define a household as ‘no college’ (‘some college’) if the male spouse has not gone to college (has some college education). Table 1 provides summary statistics. Those with some college education earn 70% more income, they have more than twice as much wealth, and they consume about 30% more.
Table 6, columns 3-4, presents the results in the quadratic model. [^45] Appendix table C.8, columns 5-8, presents the parameter estimates in the linear model. Similar to the patterns observed by wealth, we find that $\phi^{(1)}$ is higher (double) for households without college education ($0.219$ versus $0.105$). The pass-through of the average permanent shock is therefore larger among the lower educated, presumably because of the lower wealth that these households hold. $\phi^{(2)}$ is negative in both groups, suggesting that large bad permanent shocks transmit into consumption more than equal-sized good shocks. However, the magnitude of the estimate in the lower education group is over four-fold larger (and statistically significant) compared to the higher education group. Permanent shocks away from the mean are therefore way more hurtful among the lower educated. By contrast, the transmission parameters of all transitory shocks are statistically indistinguishable from zero in both groups.
@@ 6 Welfare costs of tail income risk
In the previous section, we established that income risk is negatively skewed and highly leptokurtic, and features a nonlinear pass-through to consumption. In this section, we calculate the welfare costs of such risk, particularly contrasting the implications of the linear and nonlinear consumption functions. To do so, we first draw a sequence of income shocks over the lifecycle, we then use the different consumption specifications to simulate responses to those shocks, and, finally, we use a parametric utility function to calculate household welfare. Our interest is in the costs of non-Gaussian relative to (counterfactual) Gaussian income risk.
We draw permanent and transitory income shocks for 50,000 households from mixtures of normal distributions, namely
@@
Normal mixtures offer a flexible way of modeling non-Gaussian distributions and have by now become a standard choice in quantitative income dynamics (Guvenen et al., 2021). There are five parameters per shock $(p,\mu_{1},\sigma_{1},\mu_{2},\sigma_{2})$, which are estimated using the method of simulated moments to match the first four moments of the shocks’ distributions in the PSID (column 3, table 3). We assume $\mu_{1}<0$ for identification. Appendix table D.1 reports the parameter estimates and the (excellent) model fit. In the counterfactual scenario with Gaussian risk, we fix $p=0$, $\mu_{2}=0$, and set $\sigma_{2}^{2}$ equal to the empirical variance of each shock.
@@
These reflect the model-derived linear and quadratic consumption functions (4) and (5), with $\mu^{c}_{t}=f(\Delta\mathbf{Z}_{it})$ (deterministic profile) and $\xi_{it}=0$ (mean taste shock). In practice, we let $\mu^{c}_{t}$ equal the average (non-residualized) consumption growth by year, which we estimate directly in our baseline sample. We simulate consumption over a working life of 35 years, starting at age 30. We also simulate a retirement period of 20 more years (starting at age 65), in which absence of income shocks implies $\Delta\ln C_{it}=\mu^{c}_{t}$. [^46] We fix the initial condition at age 30 ($\ln C_{it=30}$) to its mean value in the sample. In all the simulations, we restrict $\omega_{it}^{(22)}=0$ as this is never precisely estimated in any cut of the sample. Our baseline simulations use the time-invariant partial insurance parameters from table 4. Appendix table D.3 relaxes time-invariance.
We calculate the (ex-ante) welfare implied by a given consumption profile ${\cal C}_{i}=\{C_{it}\}_{t=0}^{T}$ using constant coefficient of relative risk aversion (CRRA) utility ${\cal U}({\cal C}_{i})=\sum_{t=0}^{T}\beta^{t}\frac{C_{it}^{1-\gamma}-1}{1-\gamma}$ if $\gamma\neq 1$, or logarithmic utility ${\cal U}({\cal C}_{i})=\sum_{t=0}^{T}\beta^{t}\log(C_{it})$ if $\gamma=1$. We then calculate the welfare cost of tail risk as the percentage reduction in lifetime consumption that a household who faces Gaussian income risk is willing to accept in order to avoid tail (namely, non-Gaussian) risk. This is given by $\alpha_{i}$ for a given household $i$, which, similar to the welfare calculations in Wu and Krueger (2021), satisfies the indifference equation
${\cal U}\big{(}(1-\alpha_{i}){\cal C}_{i}^{*}\big{)}={\cal U}\big{(}{\cal C}_{i}\big{)}.$
${\cal C}_{i}^{*}$ and ${\cal C}_{i}$ are alternative consumption profiles, respectively reflecting less (Gaussian) and more (non-Gaussian) income risk. We consider three alternative scenarios over which we calculate the welfare costs of non-Gaussian income risk.
In a first scenario, both ${\cal C}_{i}^{*}$ and ${\cal C}_{i}$ are simulated from the linear specification in response to Gaussian and non-Gaussian income shocks, respectively. The estimate of $\alpha_{i}$ in this case reflects the welfare cost of tail risk in the context of a linear consumption function. Tail risk is still costly in this context because the probability of very negative shocks (reflecting the long left fat tail) is much higher relative to a Gaussian density.
@@
Table 7 averages $\alpha_{i}$ over cross-sections of households defined by the magnitude of the permanent income shock they experience, in the case of $\gamma=2$. The first row presents the average cost across all households, unconditional on a given shock. This is close to zero across all three scenarios, or even negative in some cases, suggesting that households may in fact prefer tail over Gaussian risk. Inspecting the parameter estimates in table D.1, tail risk is associated with a very small probability (about 5%) that shocks come from a distribution with negative mean and large variance, and a very high probability that they come from a distribution with small positive mean but tiny variance. As such, most households will never actually experience shocks from the first distribution; their shocks will come from the latter, whose variance is an order of magnitude smaller than the empirical variance of shocks (or the variance of Gaussian risk). For many, Gaussian risk in fact implies larger variability, so households are on average willing to pay to move to the non-Gaussian environment.
@@
All households 100% -0.78 -0.86 0.65
$\min\{\zeta_{it}\}<-0.50$ 40.31% 4.74 6.31 7.71
$\min\{\zeta_{it}\}<-0.70$ 29.11% 6.28 8.53 9.90
$\min\{\zeta_{it}\}<-0.95$ 17.07% 8.43 11.74 13.06
@@
The next rows zoom in on households who experience large negative permanent shocks at least once in their lifetime. These are households who happen to draw from the first distribution in the mixture and the consequences of tail risk are clearly felt among them. Those who face a permanent shock at least three standard deviations below the mean ($\min\{\zeta_{it}\}<-0.50$; 40.31% of households experience this at least once in their lifetime) will give up 4.74% of their lifetime consumption to live in a world with Gaussian risk when such risk is transmitted linearly. This increases to 6.28% and 8.43% for shocks respectively four and five standard deviations below the mean – all shown in column 1. While these costs are already sizable, they increase further by 55% to 62% in the presence of nonlinear transmission (column 3), exceeding an eighth of lifetime consumption (13.06%) for those households who experience the worst shocks. Those estimates would be even larger if we considered the welfare cost of tail risk relative to a zero risk scenario, as, e.g., in Guvenen et al. (2024).
The welfare estimates are robust to the choice of $\gamma$ (appendix table D.2) or the use of age-specific partial insurance parameters when simulating various consumption profiles (appendix table D.3). Tail income risk consistently implies large welfare costs. The econometrician will greatly underestimate such costs if she neglects the nonlinear transmission into consumption, for which we found ample evidence in the earlier sections.
@@ 7 Conclusion
We measure the degree of consumption partial insurance to income shocks, accounting for higher-order moments of the income distribution. Such moments, in particular left skewness and excess kurtosis, are important features of unexplained income growth, documented in many recent papers. Our focus is on measuring partial insurance without taking a stance on the transmission mechanisms from income to consumption. We use recent income and consumption data from the PSID (1999–2019) and establish several new insights.
@@
Second, in estimating BPP over the recent years, we find a lower pass-through of permanent shocks (larger partial insurance) and full insurance to transitory ones. We attribute one third of the difference from BPP to the consumption imputation from the CEX, another third to the biennial nature of our data, and the remainder to time effects.
@@
Fourth, in estimating a nonlinear consumption function, we find that bad permanent shocks transmit more strongly than good ones, and the pass-through increases with the severity of the shock. By contrast, transitory shocks are fully insured in the full sample, but they are not insured among the less wealthy.
@@
In summary, the paper offers an analytical framework to measure the transmission of income shocks into consumption, allowing such transmission to vary with the sign and severity of the shock. To measure the transmission, we exploit information in higher-order income moments but we do not otherwise need to observe the entire consumption-income joint distribution (e.g., in contrast to Arellano et al., 2017). Our results reconcile the traditional partial insurance measurement literature, in which BPP is the default methodology, with both the quantitative literature (e.g., Guvenen et al., 2024) and the more empirical one based on surveys or experiments (e.g., Fuster et al., 2020), both of which document asymmetries in the pass-through of good and bad, or small and large shocks. Overall, our results suggest that tail income risk matters substantially for consumption.
@@ A.1 Taylor approximation to optimality conditions
where $\Delta c_{it+s}=\Delta\ln C_{it+s}-\theta_{it+s-1}\Delta(\mathbf{Z}_{it+s}^{\prime}\boldsymbol{\alpha})-\Xi_{it+s}$. We have split $\theta_{it+s-1}\omega_{it+s}$ into $\Xi_{it+s}$ and $\xi_{it+s}$; the first term reflects the gradient of consumption due to interest rates and impatience (discounting), the second term reflects unobserved consumption taste heterogeneity. We let demographics and time fixed effects pick up the effect of taste shifters and the gradient in the consumption path. Therefore $\Delta c_{it+s}$ can be obtained empirically from a regression of log consumption on appropriate observables (demographics, time fixed effects).

### v4

@@ 5.3.3 Results by wealth and education
We next consider four subsamples formed on the basis of household wealth and education. We use conventional definitions for ‘low’ versus ‘high’ wealth (the cutoff is median wealth in the sample) and ‘no’ versus ‘some’ college (the cutoff is 12 years of education). [^43] Wealth is defined as the sum of the value of housing, vehicles, business, investment, pension, savings, and other assets, net of mortgage and medical, student, and credit card debt. Appendix table C.1 provides summary statistics for these subsamples.
@@
Low wealth High wealth
No college Some college No college Some college
@@
Notes: The table presents the estimates of the parameters of the quadratic consumption function, allowing them to vary across wealth and education groups. The underlying income moments also vary by wealth and education (results not shown for brevity). Low (high) wealth is defined on the basis of household wealth being less (more) than median real wealth over 19991999199919992019201920192019. No versus some college is defined on the basis of the highest level of education attained by the male spouse. All columns use consumption data internally available in the PSID. Estimation is via equally weighted GMM; GMM standard errors clustered at the household level are in parentheses. Linear specification estimates appear in appendix table C.12.
Table 6 reports results from the quadratic model (see table C.12 for the linear model). Three points stand out. First, $\phi^{(1)}$ decreases monotonically as we move from group 1 (low wealth, no college) to group 4 (high wealth, some college). As $\phi^{(1)}$ measures the transmission of the average permanent shock, these groups possess gradually more insurance to average shocks; this reflects the insurance role of wealth, which increases monotonically across the groups. Second, $\phi^{(2)}$ is negative but increases (i.e., less negative) as we move across groups, with the exception of group 2 (low wealth and some college) where it equals zero. Negative shocks thus mostly transmit more strongly than positive shocks, but wealthier and more educated people have more insurance to bad shocks of given size. These people face, however, much stronger downwards income risk. This pattern is consistent with the gradient between poor and rich in table 5. Third, the transmission parameters of transitory shocks are mostly statistically insignificant, with two exceptions. In group 1 (low wealth, no college), $\psi^{(1)}$ is significant but negative, which is counterintuitive at first, yet $\psi^{(2)}$ is also negative, large, and significant. Large negative transitory shocks thus reduce consumption substantially among low wealth, low education households. In group 3 (high wealth, no college), $\psi^{(2)}$ is positive and significant, suggesting that large good shocks have a larger pass-through than smaller ones. We argued earlier that this helps reconcile the large pass-through found in quasi-experimental studies with the negligible pass-through found in studies that employ covariance restrictions. Note that the size of group 3 is by far the smallest in our sample.
5.3.4 Link with structural model
Our quasi reduced-form approach so far has enabled us to estimate an asymmetric transmission of income shocks without relying on a particular parameterization of preferences, expectations, and income distributions. Estimating a fully specified model is outside the scope of this paper, but it is useful to evaluate our findings against the few structural lifecycle models that have recently incorporated higher-order income risk.
Madera (2019) finds that earnings and nondurable consumption are correlated in the middle of the earnings distribution, but less so in the tails compared to earnings and durable consumption. Negative shocks are associated with larger pass-through rates than equal-sized positive shocks, albeit at twice the magnitude that we estimate here.
De Nardi et al. (2019) find that tail risk has large welfare costs because, among other things, it increases precautionary savings (thus also enables more insurance to a given shock). They show that the BPP method that relies on second-order moments implies a substantially higher pass-through rate than the true rate in the model.
Busch and Ludwig (2024) simulate a model with borrowing constraints. We do not model such constraints, yet a strength of our approach is that it enables us to measure partial insurance without taking a stance on the precise economic mechanisms at play. They find that the marginal propensity to consume increases with the severity of a negative permanent shock as households move towards the constraint, while it decreases with the magnitude of a positive shock. Their pass-through rates are slightly higher than ours, though their model does not feature other insurance over self-insurance.
Guvenen et al. (2024) simulate a model with retirement, stochastic lifetimes, and altruism. The consumption response varies with the size and sign of shocks; the pass-through of the average negative permanent shock is between 0.29 and 0.46, depending on age, higher than ours but lower than BPP. They show that measuring pass-through as in BPP overestimates the true rate by 1.5 times. Without non-employment shocks, agents are willing to give up almost about 22.3% of consumption to live in a world without risk.
We read these results as not too different from ours. Structural models seem to produce similar asymmetries in the transmission of income shocks, with some evidence that, as bad shocks hits, there is a limit to how much one can dissave or borrow, so those shocks end up affecting consumption by relatively more than other shocks. The magnitude of our pass-through rates seems uniformly lower than theirs (to some extent this is because these models focus on self-insurance), though other quasi reduced-form studies using the recent PSID, most notably ABB, find similar rates to ours.
@@ 6 Welfare costs of tail income risk
In this section, we calculate the welfare costs of income risk, which, as we found previously, is negatively skewed, highly leptokurtic, and features a nonlinear pass-through to consumption. We are particularly interested in contrasting the implications of linear and nonlinear transmission. To do so, we draw a sequence of income shocks over the lifecycle, we use the different consumption specifications to simulate responses to those shocks, and we use parametric utility to calculate welfare. Our interest lies in the costs of non-Gaussian relative to (counterfactual) Gaussian income risk.
Normal mixtures offer a flexible way of modeling non-Gaussian distributions and have by now become a standard choice in quantitative income dynamics (Guvenen et al., 2021, 2024). We draw permanent and transitory income shocks for 50,000 households from a mixture of two normal distributions, namely
@@
There are five parameters per shock $(p,\mu_{1},\sigma_{1},\mu_{2},\sigma_{2})$, which are estimated using the method of simulated moments to match the first four moments of shocks in the PSID (column 3, table 2). We assume $\mu_{1}<0$ for identification. Appendix table D.1 reports the parameter estimates and the (excellent) model fit. In the counterfactual scenario with Gaussian risk, we fix $p=0$, $\mu_{2}=0$, and set $\sigma_{2}^{2}$ equal to the empirical variance of each shock.
@@
These reflect the model-derived linear and quadratic consumption functions (4) and (5), with $\mu^{c}_{t}=f(\Delta\mathbf{Z}_{it})$ (deterministic profile) and $\xi_{it}=0$ (mean taste shock). In practice, we let $\mu^{c}_{t}$ equal the average (non-residualized) consumption growth by year, which we estimate in the data. We simulate consumption over a working life of 35 years, starting at age 30. In the baseline, we use the time-invariant homogeneous partial insurance parameters from table 3, but we subsequently relax time-invariance and homogeneity across households. [^44] We also simulate a retirement period of 20 years, starting at age 65, in which absence of income shocks implies $\Delta\ln C_{it}=\mu^{c}_{t}$. We fix the initial condition at age 30 ($\ln C_{it=30}$) to its mean value in the sample. In all the simulations, we restrict $\omega_{it}^{(22)}=0$ as this is never precisely estimated in any cut of the sample.
We calculate the (ex-ante) welfare implied by a given consumption profile ${\cal C}_{i}=\{C_{it}\}_{t=0}^{T}$ using constant coefficient of relative risk aversion utility ${\cal U}({\cal C}_{i})=\sum_{t=0}^{T}\beta^{t}\frac{C_{it}^{1-\gamma}-1}{1-\gamma}$ if $\gamma\neq 1$, and logarithmic utility ${\cal U}({\cal C}_{i})=\sum_{t=0}^{T}\beta^{t}\log(C_{it})$ if $\gamma=1$. We then calculate the welfare cost of tail risk as the percentage reduction in lifetime consumption that a household who faces Gaussian income risk is willing to accept in order to avoid tail (namely, non-Gaussian) risk. This is given by $\alpha_{i}$ for a given household $i$, which, similar to the welfare calculations in Wu and Krueger (2021), satisfies the indifference equation ${\cal U}\big{(}(1-\alpha_{i}){\cal C}_{i}^{*}\big{)}={\cal U}\big{(}{\cal C}_{i}\big{)}$. Here, ${\cal C}_{i}^{*}$ and ${\cal C}_{i}$ are alternative consumption profiles that reflect less (Gaussian) and more (non-Gaussian) income risk, respectively. We consider three alternative scenarios over which we calculate the welfare costs of non-Gaussian income risk.
In a first scenario, both ${\cal C}_{i}^{*}$ and ${\cal C}_{i}$ are simulated from the linear specification in response to Gaussian and non-Gaussian income shocks, respectively. The estimate of $\alpha_{i}$ in this case reflects the welfare cost of tail risk in the context of linear transmission. Tail risk is still costly in this context because the probability of very negative shocks (reflecting the long left fat tail) is much higher relative to a Gaussian density.
@@
Table 7 averages $\alpha_{i}$ over cross-sections of households defined by the magnitude of the permanent income shocks they experience, in the case of $\gamma=2$. The first row presents the average cost across all households, unconditional on a given shock. This is close to zero across all three scenarios, or even negative in some cases, suggesting that households may in fact prefer tail over Gaussian risk. Inspecting the parameter estimates in table D.1, tail risk is associated with a very small probability (about 5%) that shocks come from a distribution with negative mean and large variance, and a very high probability that they come from a distribution with small positive mean but tiny variance. As such, most households will never actually experience shocks from the first distribution; their shocks will come from the latter, whose variance is an order of magnitude smaller than the variance of Gaussian risk. For many, Gaussian risk in fact implies larger variability, so households are on average willing to pay to move to the non-Gaussian environment.
@@
All households 100% -0.81 -0.88 0.66
$\min\{\zeta_{it}\}<-0.50$ 40.31% 4.80 6.39 7.82
$\min\{\zeta_{it}\}<-0.70$ 29.11% 6.36 8.65 10.04
$\min\{\zeta_{it}\}<-0.95$ 17.07% 8.54 11.89 13.23
@@
The next rows zoom in on households who experience large negative permanent shocks at least once in their life. These are households who happen to draw from the first distribution in the mixture and the consequences of tail risk are clearly felt among them. Those who face a permanent shock at least $3\sigma$ below the mean ($\min\{\zeta_{it}\}<-0.50$; 40.31% of households experience this at least once in their lifetime) will give up 4.8% of their lifetime consumption to live in a world with Gaussian risk when such risk is transmitted linearly. This increases to 6.36% and 8.54% for shocks respectively $4\sigma$ and $5\sigma$ below the mean – all shown in column 1. These costs are already sizable, and they increase further by about 60% in the presence of nonlinear transmission (column 3), exceeding an eighth of lifetime consumption (13.23%) for those who experience the worst shocks. Those costs would be even larger if we considered a counterfactual zero risk scenario as in Guvenen et al. (2024).
These welfare calculations are ‘back-of-envelope’, as they neglect heterogeneity in income risk and consumption transmission, they treat wealth as exogenous, and they do not account for borrowing constraints. Appendix table D.2 repeats the calculations allowing for age-specific risk and consumption transmission; the costs are in the same ballpark as the baseline. Appendix table D.3 calculates the welfare costs by wealth; the least wealthy have costs almost twice as large as the baseline, while the wealthier have smaller costs, as expected. As the likelihood that a borrowing constraint binds depends on wealth or age, the age- and wealth-specific results are indirectly reflective also of such constraints. Without a structural model, we cannot allow households to switch wealth groups endogenously (e.g., in response to an income shock), yet the baseline costs lie in the middle of the wealth-specific costs, so exogenous wealth is unlikely to have caused a serious bias in our calculations. The welfare estimates are also robust to small variations in the choice of $\gamma$ (appendix table D.4).
@@ 7 Conclusion
We measure the degree of consumption partial insurance to income shocks, accounting for higher-order moments of the income distribution. Such moments, in particular left skewness and excess kurtosis, are important features of unexplained income growth, documented in many recent papers. We use recent data from the PSID and establish several new findings.
@@
Second, in estimating a linear consumption model over the recent years, we find a low pass-through of permanent shocks (large insurance) and full insurance to transitory ones. We attribute one third of the difference from BPP to the consumption imputation from the CEX, another third to the biennial nature of our data, and the remainder to time effects.
@@
Fourth, in estimating a nonlinear consumption function, we find that bad permanent shocks transmit more strongly than good ones, and the pass-through increases with the severity of the shock. By contrast, transitory shocks are fully insured in the full sample, but they are not insured among the less wealthy. The nature of income risk and consumption transmission vary by age and the household’s position in the income distribution.
@@
In summary, the paper offers a straightforward analytical framework to measure the transmission of income shocks into consumption, allowing such transmission to vary with the sign and size of the shock. To do so, we exploit higher-order income moments but we do not otherwise need to observe the entire consumption-income joint distribution. Our results consistently point to asymmetries in the pass-through of good and bad shocks, or by age and income level, suggesting that tail income risk matters substantially for consumption.
@@ A.1 Taylor approximation to optimality conditions
where $\Delta c_{it+s}=\Delta\ln C_{it+s}-\theta_{it+s-1}\Delta(\mathbf{Z}_{it+s}^{\prime}\boldsymbol{\alpha})-\Xi_{it+s}$. We have split $\theta_{it+s-1}\omega_{it+s}$ into $\Xi_{it+s}$ and $\xi_{it+s}$; the first term reflects the gradient of consumption due to interest rates and impatience (discounting), the second term reflects unobserved consumption taste heterogeneity. We let demographics and time fixed effects pick up the effect of taste shifters and the gradient in the consumption path. Therefore $\Delta c_{it+s}$ can be obtained empirically from a regression of consumption growth on appropriate observables (demographics, time fixed effects).